# Rethinking Multi-domain Generalization with A General Learning Objective

### v1

@@ 7.1 Derivation details of PUB
Details for Eq. 6. Taking into account $\mathcal{O}$, similar to [9], we have the upper bound for GJSD as:
@@
Details for Eq. 7. The above bound can be further reformed as:
@@
where is GReg2.
Derivation details of Eq. 10. For a Gaussian distribution $\mathcal{N}(x;\mu,\Sigma)$ with $D$ dimension, its entropy is:
@@
Then Eq. 10 equals:
@@
$+H(\mathbb{E}[P({\psi({\mathbf{Y}}_{n})\mid\phi({\mathbf{X}}_{n})})])\;]$ (40)
@@
$+H(\mathbb{E}[P({\psi({\mathbf{Y}}_{n})\mid\phi({\mathbf{X}}_{n})})]),$ (41)
@@
Since the entropy is non-negative and the constants can be omitted, Eq. 46 is equivalent to:
@@
By grouping Eq. 47, we have that for Aim 1 it minimizes $\min_{\phi}H(\mathbb{E}[P({\phi({\mathbf{X}}_{n}),{\mathbf{Y}}_{n}})])$, and for Aim 2 it minimizes $H(P({\phi({\mathbf{X}})\mid{\mathbf{Y}}}))-H(\mathbb{E}[P({\phi({\mathbf{X}})\mid{\mathbf{Y}}})]))+H(P(\phi({\mathbf{X}})))$.
@@ 11.1 Other findings and Analysis
What makes a better $\mathcal{O}$. As demonstrated in Eq. 7, $\mathcal{O}$ plays a crucial role in PUB by anchoring a space where the relationship between ${\mathbf{X}}$ and ${\mathbf{Y}}$ is preserved. Ideally, having one $\mathcal{O}$ that provides general representations for all seen and unseen domains leads to the best results, one finding supported by MIRO and SIMPLE. However, even though SIMPLE++ combines 283 pre-trained models, achieving the ‘perfect’ $\mathcal{O}$ remains unattained. Therefore, this paper primarily discusses how our proposed objectives can improve the model performance when a fixed $\mathcal{O}$ is provided.

### v2

@@ 7.1 Derivation details of PUB
Details for Eq. LABEL:eq:bound2. Taking into account $\mathcal{O}$, similar to [9], we have the upper bound for GJSD as:
@@
Details for Eq. LABEL:eq:PUB. The above bound can be further reformed as:
@@
which is GReg2.
Derivation details of Eq. LABEL:eq:cfs_detial. For a Gaussian distribution $\mathcal{N}(x;\mu,\Sigma)$ with $D$ dimension, its entropy is:
@@
Then Eq. LABEL:eq:cfs_detial equals:
@@
$\underbrace{+H(\mathbb{E}[P({\psi({\mathbf{Y}}_{n})\mid\phi({\mathbf{X}}_{n})})])}_{\text{Minimized.}}\;]$ (40)
@@
$\underbrace{+H(\mathbb{E}[P({\psi({\mathbf{Y}}_{n})\mid\phi({\mathbf{X}}_{n})})])}_{\text{Remains.}},$ (41)
@@
Since the entropy is non-negative and the constants can be omitted, Eq. LABEL:eq:MDA is equivalent to:
@@
By grouping Eq. LABEL:eq:MDA2, we have that for Aim 1 it minimizes $\min_{\phi}H(\mathbb{E}[P({\phi({\mathbf{X}}_{n}),{\mathbf{Y}}_{n}})])$, and for Aim 2 it minimizes $H(P({\phi({\mathbf{X}})\mid{\mathbf{Y}}}))-H(\mathbb{E}[P({\phi({\mathbf{X}})\mid{\mathbf{Y}}})]))+H(P(\phi({\mathbf{X}})))$.
@@ 11 More results
Visualization of toy experiments: Please see the visualization of toy experiments in Figure 4.
@@ 11.1 Other findings and Analysis
What makes a better $\mathcal{O}$. As demonstrated in Eq. LABEL:eq:PUB, $\mathcal{O}$ plays a crucial role in PUB by anchoring a space where the relationship between ${\mathbf{X}}$ and ${\mathbf{Y}}$ is preserved. Ideally, having one $\mathcal{O}$ that provides general representations for all seen and unseen domains leads to the best results, one finding supported by MIRO and SIMPLE. However, even though SIMPLE++ combines 283 pre-trained models, achieving the ‘perfect’ $\mathcal{O}$ remains unattained. Therefore, this paper primarily discusses how our proposed objectives can improve the model performance when a fixed $\mathcal{O}$ is provided.